OpenStudy (anonymous):

a point moves so that the difference between its distance from (0, 5) and (0, -5) is 8. find the equation of the locus.

OpenStudy (anonymous):

Sounds like a hyperbola.

OpenStudy (anonymous):

yeah hyperbola again..

OpenStudy (anonymous):

They points they gave you are the foci.

OpenStudy (anonymous):

wio, how did your SS jump to 90? weezz

OpenStudy (anonymous):

Well, I was at 89 before.

OpenStudy (anonymous):

I know, but still! Once upon a time you were at 70 or something around that

OpenStudy (anonymous):

You lucky fella! [:

OpenStudy (anonymous):

Nonsense, this past week I've been in the upper 80's!

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

yw! o.o

OpenStudy (anonymous):

\[ \sqrt{(5-b)^2}=\sqrt{(-5-b)^2} \]

OpenStudy (anonymous):

Doing this can at least find one of the vertices! @silverxx

OpenStudy (anonymous):

We know there is a vertex on \((0,b)\).

OpenStudy (anonymous):

can we use also distance formula right?

OpenStudy (anonymous):

That is what I just did up there.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[ \sqrt{(5-b)^2}=\sqrt{(-5-b)^2} \] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

ah okay okay

OpenStudy (anonymous):

\[ (25-10b+b^2)=(25+10b+b^2)+8 \]

OpenStudy (anonymous):

I should have wrote:\[ \sqrt{(5-b)^2}-\sqrt{(-5-b)^2}=8 \]

OpenStudy (anonymous):

b is 2/5?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

and how about a?

OpenStudy (anonymous):

the answer is 9y^2 - 16x^2 = 144

OpenStudy (anonymous):

stated on the book

OpenStudy (anonymous):

Don't cheat!!!

OpenStudy (anonymous):

@ⒶArchie☁✪ Do you know how to go from here?

OpenStudy (anonymous):

no no. i mean this question was one of the questions came up during our quiz and i need to know the steps on how to achieve the right answer..

OpenStudy (anonymous):

professor gets the questions on the book..

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

Okay, the next step is to find latus rectum.

OpenStudy (anonymous):